OpenStudy (anonymous):

Solve the system algebraically y=x^2-x-6 and y=2x-2

OpenStudy (anonymous):

set the y's equal: x^2-x-6 = 2x-2 x^2 - 3x - 4 = 0 can you solve this from here? HINT: factor or QF

OpenStudy (anonymous):

i did I got the points (4,6) and (-1,-4) but the 2nd one doesnt work

OpenStudy (anonymous):

the second one, (-1, -4) does work ...

OpenStudy (anonymous):

no way.

OpenStudy (anonymous):

no ot for y=x^2-x-6

OpenStudy (anonymous):

(-1)^2 -(-1)-6 = 1 + 1 - 6 = 2 - 6 = -4

OpenStudy (anonymous):

you a genius so its both

OpenStudy (anonymous):

righyt....